OpenStudy (anonymous):

Use the given circumference to approximate the surface area of the spherical object. Round the answer to the nearest square centimeter. c = 27 cm @Directrix

Directrix (directrix):

What kind of object is shown. Is it a complete sphere?

OpenStudy (anonymous):

its a bowling ball

Directrix (directrix):

First, use the circumference formula to find the radius of the bowling bowl.

Directrix (directrix):

C = 2* pi*r 27 = 2 * 3.14 * r Solve for r.

Directrix (directrix):

Did you get r = 4.3 (rounded)

OpenStudy (anonymous):

i dont know how to get r

Directrix (directrix):

Back to the surface area of the sphere: A = 4 * π * r² A = 4 * π * (4.3)²

OpenStudy (anonymous):

232

Directrix (directrix):

Yes, 232 to the nearest square cm

Directrix (directrix):

next question